OpenStudy (anonymous):

Absolute Value Equations Does this problem have a solution that is equal to one another? |x + 4| = 3x + 2

OpenStudy (anonymous):

there is only one solution here, although sometimes there are two

OpenStudy (anonymous):

i got 2.5 = 2.5 but its supposed to be 2.5 = - 2.5 what did i do wrong?

OpenStudy (anonymous):

first of all \(1\) is a solution "by instpecton" since it is pretty clear that \[|1+4|=3+2\]

OpenStudy (anonymous):

\(2.5\) is not a solution

OpenStudy (anonymous):

through my experience with absolute value always one answer is positive and one answer is negative but they're the same number

OpenStudy (anonymous):

not true here however

OpenStudy (anonymous):

im talkin about the negative side oops sorry forgot to mention that

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i have for the positive x = 1 and 5 = 5 and for the negative i have x = 1.5 and 2.5 = 2.5

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

|x +4| = 3x + 2 is the positive and |x + 4| = -3x - 2 is the negative

OpenStudy (anonymous):

They will have the same magnitude, but opposing signs you mean.

OpenStudy (anonymous):

if \(x>-4\) then \(|x+4|=x+4\) and you can solve \[x+4=3x+2\] easly

OpenStudy (anonymous):

you get \[x=1\]

OpenStudy (anonymous):

yea but the other side is negative for the second half of the problem

OpenStudy (anonymous):

and its -1.5

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

and i plug it in

OpenStudy (anonymous):

if \(x<-4\) then \[|x+4|=-x-4\] and lets see what happens when we solve \[-x-4=3x+2\]

OpenStudy (anonymous):

we get \[4x=-6\] so \[x=-\frac{3}{2}=-1.5\] HOWEVER

OpenStudy (anonymous):

we were making the assumption that \(x<-4\) and so \(x\) CANNOT be \(-1.5\) since \(-1.5>-4\)

OpenStudy (anonymous):

conclude that there is only one solution \(x=1\) alternatively graph the line \(y=3x+2\) and the V shaped \(y=|x+4|\) and see that they intersect in only one place

OpenStudy (anonymous):

wait you're losing me so x cant be greater than -4?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

thats what she said

OpenStudy (anonymous):

it is clear that if you solve \[x+4=2x+3\] you get \(x=1\) right actually she said "if you want to make me happy you gotta make it snappy"

OpenStudy (anonymous):

yea i get that part

OpenStudy (anonymous):

https://www.youtube.com/watch?v=wiPizq1OEG8

OpenStudy (anonymous):

ok so then the other thing to solve is \[-x-4=3x+2\]

OpenStudy (anonymous):

but \(|x+4|=-x-4\)

OpenStudy (anonymous):

ONLY IF \(x<-4\) otherwise \(|x+4|=x+4\)

OpenStudy (anonymous):

so when you solve \[-x-4=3x+2\] and get some number LARGER than \(-4\) it is not a real solution

OpenStudy (anonymous):

wait u lost me with -x-4 = 3x+2 why did you do that?

OpenStudy (anonymous):

because \(|x+4|\) is either \(x+4\) or its negative \(-x-4\) what two equation would you solve?

OpenStudy (anonymous):

hey is that a rule? if the number is larger than the number all the way to the right that its a no solution?

OpenStudy (anonymous):

i don't know if it is a rule, but it is true

OpenStudy (anonymous):

i bet you were going to solve \[x+4=-3x-2\] right?

OpenStudy (anonymous):

yea!!

OpenStudy (anonymous):

why isnt that right???

OpenStudy (anonymous):

well convince yourself that that is the same as solving \[-x-4=3x+2\]

OpenStudy (anonymous):

just multiply both sides by \(-1\)

OpenStudy (anonymous):

so go ahead, solve \[x+4=-3x-2\] and when you check your answer you will see it does not work !!

OpenStudy (anonymous):

yea its not supposed to work the guy in the GRE video got 2.5 and a -2.5 thats the right answer

OpenStudy (anonymous):

im tryna figure out how he got the 2nd negative cuz i keep gettin 2.5 for both

OpenStudy (anonymous):

there is no \(2.5\) in it i see what you mean though once you solve, you get \[x=-1.5\] then substitute in to the original equation

OpenStudy (anonymous):

yea!! thats what i did and it came out to be 2.5 for both sides

OpenStudy (anonymous):

but its supposed to be -2.5 on the right side .... i don't kno how he got that

OpenStudy (anonymous):

because you did not substitute in to the ORIGINAL equation \[|x + 4| = 3x + 2\] you substituted in to \[x+4=-3x-2\]

OpenStudy (anonymous):

it is a solution to the second equation, but not to the first

OpenStudy (anonymous):

wait i thought you were supposed to substitute the what you got from the negative to the negative problem?

OpenStudy (anonymous):

you have to plug in both x=1 and x=-1.5 to the first equation not the second?

OpenStudy (anonymous):

yes, you are solving \[|x + 4| = 3x + 2\] right ? that is what you have to check

OpenStudy (anonymous):

forgot how much i liked george clinton

OpenStudy (anonymous):

oh wth............. i did not know that.... i see it now

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

now i am going to go listen to more funk later

OpenStudy (anonymous):

nice !!